a ball of mass 0.2 kg rests on a vertical post of length 5m.


a bullet of mass 0.01 kgtravelling with a vel v m/s in horizonal direction,hits center of ball


after collision bullet and ball travell seperately


ball hits ground at dist of 20m from post and bullet at 100m from post


then find initial velocity of the bullet?


A)250


B)250*(2)^1/2


C)400


D)500


Grade:11

1 Answers

Pranav -
10 years ago

D- The question is about conservation of momentum

Assume the ball and the bullet after collision travelled their respective distances in 1sec, which makes there speed 20m/s and 100m/s.

You can calculate the momentum for each of them using the formule p=mv

0.2*20=4kgm/s

0.01*100=1kgm/s

Add the momentums (4+1=5) and that is the transferred by the bullet.

And now you can find out the velocity because you have the mass and momentum of the bullet at the start.

p=mv

v=5/0.01

v=500m/s

Thank you